# Berkovich Motives

### v1

@@ 10.1. Mixed Tate motives
For any strictly totally disconnected $X$, we let
$\mathcal{D}_{MT}(X)\subset\mathcal{D}_{\mathrm{mot}}(X)$
be the full subcategory generated by all Tate twists $\mathbb{Z}(n)$, $n\in\mathbb{Z}$; this is the “mixed Tate” subcategory. The functor
$X\mapsto\mathcal{D}_{MT}(X)$
from strictly totally disconnected $X$ to compactly generated presentable stable $\infty$-categories (and functors preserving compact objects and colimits) preserves finite products and filtered colimits, and hence extends uniquely to an arc-sheaf, defining $\mathcal{D}_{MT}(X)\subset\mathcal{D}_{\mathrm{mot}}(X)$ for general $X$ via descent. Here, note that the $\infty$-category $\mathrm{Pr}^{\omega}\cong\mathrm{Cat}_{\infty}^{\mathrm{idem}}$ of compactly generated presentable $\infty$-categories and functors preserving compact objects and filtered colimits (equivalently, the $\infty$-category of small idempotent-complete $\infty$-categories) is itself compactly generated.
@@
An object $M\in\mathcal{D}_{\mathrm{mot}}(X)$ lies in $\mathcal{D}_{MT}(X)$ if and only if its fibre at all geometric points $\mathcal{M}_{\mathrm{arc}}(C(x))\to X$ lies in $\mathcal{D}_{MT}(C(x))$.
@@
By descent, we can assume that $X$ is strictly totally disconnected. The inclusion $\mathcal{D}_{\mathrm{MT}}(X)\subset\mathcal{D}_{\mathrm{mot}}(X)$ has a right adjoint $R$ commuting with all colimits (as the inclusion preserves compact objects). The map $R(M)\to M$ in $\mathcal{D}_{\mathrm{mot}}(X)$ becomes an isomorphism at all stalks by the assumption and the commutation of $R$ with colimits. Thus, it is an isomorphism. ∎
@@ 11. $\mathcal{D}_{\mathrm{mot}}(k)$
Using the results of the previous section, we can now also understand the structure of $\mathcal{D}_{\mathrm{mot}}(k)$, and in fact compare it to Voevodsky’s theory, for any (discrete) field $k$, considered as a Banach ring with the trivial norm. We consider the case that $k$ is algebraically closed; one can reduce the general case to this case by Galois descent.
@@
is, after rationalization, given by the weight $j$ Adams piece of $K(X)$; while after profinite completion (prime to the characteristic $p$ of $k$) it is given by (pro)étale cohomology of $X$ with coefficients in ^$\hat{\mathbb{Z}}^{p}(j)$.
@@
Let $K=k((T))_{1/2}$ be the Laurent series on a free variable $T$ with exact norm $\tfrac{1}{2}$, with $f:\mathcal{M}_{\mathrm{arc}}(K)\to\mathcal{M}_{\mathrm{arc}}(k)$. This is a closed subset of $g:\mathbb{G}_{m,k}\to\mathcal{M}_{\mathrm{arc}}(k)$, and the map $g_{\ast}\mathbb{Z}\to f_{\ast}\mathbb{Z}$ is an isomorphism; the cone of $\mathbb{Z}\to f_{\ast}\mathbb{Z}$ is isomorphic to (1)[$\mathbb{Z}(-1)[-1]$. In particular, $f_{\ast}\mathbb{Z}$ is dualizable. We claim that its dual $(f_{\ast}\mathbb{Z})^{\vee}$ is representing the functor $M\mapsto\mathrm{Hom}(1,f^{\ast}M)$. Indeed,
@@
In particular, the left adjoint to $f^{\ast}$ exists on the unit, with value $(f_{\ast}\mathbb{Z})^{\vee}$. As $\mathcal{D}_{\mathrm{mot}}(K)$ is generated by the image of $f^{\ast}$ and Artin motives (which can be handled by passing to finite extensions of $K$, which are also Laurent series fields), it follows that the left adjoint $f_{\sharp}$ exists in general.
It also follows that $f_{\sharp}\mathbb{Z}=(f_{\ast}\mathbb{Z})^{\vee}$ is a compact object. But $f_{\ast}\mathbb{Z}\cong g_{\ast}\mathbb{Z}$ has $\mathbb{Z}$ as a summand, using say the unit section $1\in\mathbb{G}_{m,k}$. Thus $\mathbb{Z}$ is also compact. Moreover, $\mathcal{D}_{\mathrm{mot}}(k)$ is generated by $f_{\sharp}\mathcal{D}_{\mathrm{mot}}(K)$, which in turn is generated by
@@
which can also be written as the pushout
[…]
@@
But writing $\mathcal{O}_{C}$ as a completed filtered colimit of finite extensions of $k[[T]]$, and using excision on regular schemes, the term $K(X_{\mathcal{O}_{C}}\ \mathrm{on}\ X)_{\mathbb{Q}}$ is isomorphic to $K(X)_{\mathbb{Q}}$; with Adams weight spaces shifted by $1$. This extra piece matches the second summand of $f_{\sharp}\mathbb{Z}$. ∎
@@
is an isomorphism, by direct computation. But $g_{\ast}\mathbb{Z}$ has a natural map back to $\mathbb{Z}$ given by $1\in\tilde{\mathbb{G}}_{m,k}$.
@@
The map $f_{\ast}\mathbb{Z}\cong g_{\ast}\mathbb{Z}\to\mathbb{Z}$ is descendable. In particular, $\Psi$ is conservative, and we may regard $\mathcal{D}_{\mathrm{mot}}(C)$ as a “neutral Tannakian stable $\infty$-category over $\mathcal{D}_{\mathrm{mot}}(k)$”.
@@
The fibre of $g_{\ast}\mathbb{Z}\to\mathbb{Z}$ is isomorphic to (1)[$\mathbb{Q}(-1)[-1]$ and the $g_{\ast}\mathbb{Z}$-module structure on it factors over $g_{\ast}\mathbb{Z}\to\mathbb{Z}$: Indeed, this can be checked rationally, and rationally $g_{\ast}\mathbb{Q}$ is the free $E_{\infty}$-algebra on (1)[$\mathbb{Q}(-1)[-1]$. Thus, $g_{\ast}\mathbb{Z}$ is filtered by modules inflated along $g_{\ast}\mathbb{Z}\to\mathbb{Z}$, which implies descendability. ∎
@@
Note that $\mathcal{D}_{\mathrm{mot}}(k)$ is naturally linear over $\mathcal{D}_{\mathrm{qc}}(X)$. Indeed, the map to $\ast/\mathbb{G}_{m}$ is given by the Tate twist, while the lift to the ^$\hat{\mathbb{Z}}^{p}(1)$-torsor over it is given by the isomorphisms
@@
The functor $\mathcal{D}_{\mathrm{qc}}(X)\to\mathcal{D}_{\mathrm{mot}}(k)$ is fully faithful, with essential image being the subcategory generated under colimits, shifts, and retracts by $\mathbb{Z}(n)$ for $n\in\mathbb{Z}$. This agrees with $\mathcal{D}_{MT}(k)\subset\mathcal{D}_{\mathrm{mot}}(k)$.
@@
Note that the unit in $\mathcal{D}_{\mathrm{qc}}(X)$ is compact, and $\mathcal{D}_{\mathrm{qc}}(X)$ is generated by the tensor powers of the line bundle $L$; this follows from $X$ being affine over $\ast/\mathbb{G}_{m}$. To prove the proposition, it remains to compare homomorphisms between generating objects, and this can be split up into the case of torsion coefficients, and rational coefficients. The case of modulo $n$ coefficients is clear as both categories become $\mathcal{D}(\mathbb{Z}/n\mathbb{Z})$, while the rational case follows from the previous proposition.
It is clear that the subcategory generated by all $\mathbb{Z}(n)$ is contained in $\mathcal{D}_{MT}(k)$. For the converse, pull back to $\mathcal{D}_{MT}(C)$ and use that every object is a retract of push-pull along $\mathcal{M}_{\mathrm{arc}}(C)\to\mathcal{M}_{\mathrm{arc}}(k)$. ∎
We note that it follows that $\mathcal{D}_{MT}(k)$ has a natural $t$-structure coming from the equivalence with $\mathcal{D}_{\mathrm{qc}}(X)$. This is the “motivic $t$-structure”, whose connective part is the one generated under colimits by $\mathbb{Z}(n)$, $n\in\mathbb{Z}$; after $\ell$-adic completion, it gives the usual $t$-structure on $\ell$-adic sheaves.
@@
induced by the map $f_{\ast}\mathbb{Z}\cong g_{\ast}\mathbb{Z}\to\mathbb{Z}$. It suffices to see that $f_{\ast}\mathbb{Z}\to\mathbb{Z}$ is coconnectively faithfully flat in $\mathcal{D}_{MT}(k)\cong\mathcal{D}_{\mathrm{qc}}(X)$, i.e. tensoring preserves coconnective modules, and is conservative on them. Covering $X$ by $\mathrm{Spec}(\mathbb{Z}[\tfrac{1}{p}])$, we have to see that the map from the free $E_{\infty}$-$\mathbb{Z}[\tfrac{1}{p}]$-algebra on $\mathbb{Q}[-1]$ back to $\mathbb{Z}[\tfrac{1}{p}]$ is coconnectively faithfully flat. This is a simple computation. ∎
Let $X_{C}\to X$ be the algebraic stack parametrizing, in addition to the line bundle $L/R$ with compatible isomorphisms $L/n\cong\mu_{n}(k)\otimes_{\mathbb{Z}}R$, a map
@@
that after reduction modulo $n$ yields the natural map
@@
which after pullback to $\mathcal{O}_{C}^{\times}/(1+C_{<1})\otimes_{\mathbb{Z}}R\cong L\otimes\mathbb{Q}/\mathbb{Z}[\tfrac{1}{p}]$ is identified with
@@
As $C^{\times}/\mathcal{O}_{C}^{\times}\cong\mathbb{Q}$, the map $X_{C}\to X$ is a torsor under $\mathrm{Hom}(\mathbb{Q},L)$. This is an affine group scheme, and in fact $h:X_{C}\to X$ is a relative affine stack, so
$\mathcal{D}_{\mathrm{qc}}(X_{C})\cong\mathrm{Mod}_{h_{\ast}\mathcal{O}}\mathcal{D}_{\mathrm{qc}}(X).$
Moreover, $h_{\ast}\mathcal{O}$ is a complex which in degree 00 is given by $\mathcal{O}$, and in degree $1$ by $L^{\vee}\otimes\mathbb{Q}$, and 00 otherwise.
The identification $\mathbb{Z}(1)(C)\cong C^{\times}/(1+C_{<1})[-1]$ yields a natural $\mathcal{D}_{\mathrm{qc}}(X)\cong\mathcal{D}_{MT}(k)$-linear symmetric monoidal functor
$\mathcal{D}_{\mathrm{qc}}(X_{C})\to\mathcal{D}_{MT}(C).$
@@
The functor $\mathcal{D}_{\mathrm{qc}}(X_{C})\to\mathcal{D}_{\mathrm{MT}}(C)$ is an equivalence. Consequently,
$\mathcal{D}_{\mathrm{mot}}(C)\cong\mathcal{D}_{\mathrm{mot}}(k)\otimes_{\mathcal{D}_{\mathrm{qc}}(X)}\mathcal{D}_{\mathrm{qc}}(X_{C}).$
@@
By our preparations, it suffices to see that $h_{\ast}\mathcal{O}\cong f_{\ast}\mathbb{Z}$ as $E_{\infty}$-algebras under the equivalence $\mathcal{D}_{\mathrm{qc}}(X)\cong\mathcal{D}_{MT}(k)$. The constructions give a natural map, and we know both sides explicitly as complexes: They are $\mathbb{Z}[\tfrac{1}{p}]\oplus\mathbb{Q}(-1)[-1]$. One only has to see that the map (1)[$\mathbb{Q}(-1)[-1]\to\mathbb{Q}(-1)[-1]$ is an isomorphism. But such endomorphisms are given by multiplication by some rational number, and it is easy to see that it must be nonzero. ∎
@@
and consequently a section of $X_{C}\to X$; this is responsible for the nearby cycles functor

### v2

@@ 10.1. Mixed Tate motives
For any small arc-stack $X$, we define a subcategory of mixed Tate motives $\mathcal{D}_{MT}(X)\subset\mathcal{D}_{\mathrm{mot}}(X)$, as follows. (This might more properly be called mixed Artin–Tate, as we check the condition only at geometric points.)
Definition 10.7.
An object $M\in\mathcal{D}_{\mathrm{mot}}(X)$ lies in $\mathcal{D}_{MT}(X)$ if its fibre at all geometric points $\mathcal{M}_{\mathrm{arc}}(C)\to X$, $C$ a nondiscrete algebraically closed Banach field, lies in the subcategory generated by all Tate twists $\mathbb{Z}(n)$, $n\in\mathbb{Z}$.
@@
The functor $X\mapsto\mathcal{D}_{MT}(X)$ defines an arc-sheaf of (symmetric monoidal presentable stable) $\infty$-categories.
@@
It suffices to see that if an object $M\in\mathcal{D}_{\mathrm{mot}}(X)$ lies in $\mathcal{D}_{MT}(X)$ arc-locally, then it lies in $\mathcal{D}_{MT}(X)$. From the definition, it suffices to check this when $X=\mathcal{M}_{\mathrm{arc}}(C)$ and for an arc-cover by $X^{\prime}=\mathcal{M}_{\mathrm{arc}}(C^{\prime})$. This is nontrivial only when $C$ is nonarchimedean. Take any $M\in\mathcal{D}_{\mathrm{mot}}(X)$. As $\mathcal{D}_{\mathrm{mot}}(X)$ is compactly generated, we can write $M=\mathrm{colim}_{i}M_{i}$ as a filtered colimit of compact $M_{i}$. Assume that $M|_{C^{\prime}}$ is mixed-Tate. Then there is another presentation $M|_{C^{\prime}}=\mathrm{colim}_{j}M^{\prime}_{j}$ where all $M^{\prime}_{j}$ are compact and mixed-Tate. The two presentations $\mathrm{colim}_{i}M_{i}|_{C^{\prime}}=M|_{C^{\prime}}=\mathrm{colim}_{j}M^{\prime}_{j}$ must be Ind-isomorphic. It follows that for any $i$, there is some $i^{\prime}\geq i$ so that the map $M_{i}|_{C^{\prime}}\to M_{i^{\prime}}|_{C^{\prime}}$ factors over a mixed-Tate object. It suffices to see that in this case also $M_{i}\to M_{i^{\prime}}$ factors over a mixed-Tate object. To see this, write $C^{\prime}$ as a completed filtered colimit of algebraically closed Banach fields $C_{k}$ that have finite topological transcendence degree over $C$. Both the mixed-Tate object and the factorization of $M_{i}|_{C^{\prime}}\to M_{i^{\prime}}|_{C^{\prime}}$ over it are then already defined over some $C_{k}$; so we may assume that $C^{\prime}$ has finite topological transcendence degree. We can then write $C^{\prime}$ as a completed filtered colimit of smooth $C$-algebras topologically of finite type $B_{k}$, of uniformly bounded dimension. Then again, the mixed-Tate object and the factorization $M_{i}|_{C^{\prime}}\to M_{i^{\prime}}|_{C^{\prime}}$ over it are already defined over some $B_{k}$. But $B_{k}$ admits a section $B_{k}\to C$, thus yielding the factorization already over $C$. ∎
For strictly totally disconnected $X$, the point-wise condition in the definition becomes global.
Proposition 10.9.
For strictly totally disconnected $X$, the subcategory
$\mathcal{D}_{MT}(X)\subset\mathcal{D}_{\mathrm{mot}}(X)$
is the full subcategory generated by all Tate twists $\mathbb{Z}(n)$, $n\in\mathbb{Z}$.
Proof.
Let $\mathcal{D}_{MT}^{\prime}(X)\subset\mathcal{D}_{\mathrm{mot}}(X)$ be the subcategory generated by all Tate twists. The inclusion $\mathcal{D}_{\mathrm{MT}}^{\prime}(X)\subset\mathcal{D}_{\mathrm{mot}}(X)$ has a right adjoint $R$ commuting with all colimits (as the inclusion preserves compact objects). Assume that $M\in\mathcal{D}_{MT}(X)$. The map $R(M)\to M$ in $\mathcal{D}_{\mathrm{mot}}(X)$ becomes an isomorphism at all stalks by the assumption and the commutation of $R$ with colimits. Thus, it is an isomorphism. Therefore $\mathcal{D}_{MT}(X)\subset\mathcal{D}_{MT}^{\prime}(X)$, and the other inclusion is clear. ∎
@@ 11. $\mathcal{D}_{\mathrm{mot}}(k)$
Using the results of the previous section, we can now also understand the structure of $\mathcal{D}_{\mathrm{mot}}(k)$, and in fact compare it to Voevodsky’s theory, for any (discrete) field $k$, considered as a Banach ring with the trivial norm. We consider the case that $k$ is algebraically closed; one can reduce the general case to this case by Galois descent (but may have to be careful with descent versus hyperdescent questions, depending on the precise definitions used).
@@
is, after rationalization, given by the weight $j$ Adams piece of $K(X)$; while after profinite completion (prime to the characteristic $p$ of $k$) it is given by proétale cohomology of $X$ with coefficients in $\hat{\mathbb{Z}}^{p}(j)$.
@@
Let $K=k((T))_{1/2}$ be the Laurent series on a free variable $T$ with exact norm $\tfrac{1}{2}$, with $f:\mathcal{M}_{\mathrm{arc}}(K)\to\mathcal{M}_{\mathrm{arc}}(k)$. This is a closed subset of $g:\mathbb{G}_{m,k}\to\mathcal{M}_{\mathrm{arc}}(k)$, and the map $g_{\ast}\mathbb{Z}\to f_{\ast}\mathbb{Z}$ is an isomorphism (use Proposition 5.19, pass to duals, and pass to limits resp. colimits over increasing resp. shrinking annuli, using finitaryness). The cone of $\mathbb{Z}\to f_{\ast}\mathbb{Z}$ is isomorphic to $\mathbb{Z}(-1)[-1]$. In particular, $f_{\ast}\mathbb{Z}$ is dualizable. We claim that its dual $(f_{\ast}\mathbb{Z})^{\vee}$ is representing the functor $M\mapsto\mathrm{Hom}(1,f^{\ast}M)$. Indeed,
@@
In particular, the left adjoint to $f^{\ast}$ exists on the unit, with value $(f_{\ast}\mathbb{Z})^{\vee}$. By Proposition 10.1, $\mathcal{D}_{\mathrm{mot}}(K)$ is generated by the image of $f^{\ast}$ and Artin motives (which can be handled by passing to finite extensions of $K$, which are also Laurent series fields), it follows that the left adjoint $f_{\sharp}$ exists in general.
It also follows that $f_{\sharp}\mathbb{Z}=(f_{\ast}\mathbb{Z})^{\vee}$ is a compact object. But $f_{\ast}\mathbb{Z}\cong g_{\ast}\mathbb{Z}$ has $\mathbb{Z}$ as a summand, using say the unit section $1\in\mathbb{G}_{m,k}$. Thus $\mathbb{Z}$ is also compact. Moreover, $\mathcal{D}_{\mathrm{mot}}(k)$ is generated by $f_{\sharp}\mathcal{D}_{\mathrm{mot}}(K)$ as $f^{\ast}$ is conservative, which in turn is generated by
@@
(where we write $K(X_{\mathfrak{m}_{C}})$ for the fiber of $K(X_{\mathcal{O}_{C}})\to K(X)$), which can also be written as the pushout
${K(X_{\mathcal{O}_{C}})_{\mathbb{Q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K(X_{C})_{\mathbb{Q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${K(X)_{\mathbb{Q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\overline{K}(X_{C})_{\mathbb{Q}}}$
@@
But writing $\mathcal{O}_{C}$ as a completed filtered colimit of finite extensions of $k[[T]]$, and using excision on regular schemes, the term $K(X_{\mathcal{O}_{C}}\ \mathrm{on}\ X)_{\mathbb{Q}}$ is isomorphic to $K(X)_{\mathbb{Q}}$; with Adams weight spaces shifted by $1$. This extra piece matches the second summand of $f_{\sharp}\mathbb{Z}$ (argue by induction on Adams weight). ∎
@@
is an isomorphism, by direct computation. (Indeed, by adjunction both sides have compatible maps from $\mathbb{Z}$, and the cofibre is $\mathbb{Q}(-1)[-1]$. This follows by passing to colimits in the description at finite levels; for $\tilde{\mathbb{G}}_{m}$ this has to be combined with a limit over increasing annuli, noting that all transition maps in the limit are isomorphisms. The induced endomorphism of $\mathbb{Q}(-1)[-1]$ is nonzero, by comparison to any finite level, so must be an isomorphism.) But $g_{\ast}\mathbb{Z}$ has a natural map back to $\mathbb{Z}$ given by $1\in\tilde{\mathbb{G}}_{m,k}$.
@@
The map $f_{\ast}\mathbb{Z}\cong g_{\ast}\mathbb{Z}\to\mathbb{Z}$ is descendable in the sense of [Mat16, Definition 3.18]. In particular, $\Psi$ is conservative, and we may regard $\mathcal{D}_{\mathrm{mot}}(C)$ as a “neutral Tannakian stable $\infty$-category over $\mathcal{D}_{\mathrm{mot}}(k)$”.
@@
The fibre of $g_{\ast}\mathbb{Z}\to\mathbb{Z}$ is isomorphic to $\mathbb{Q}(-1)[-1]$ and the $g_{\ast}\mathbb{Z}$-module structure on it factors over $g_{\ast}\mathbb{Z}\to\mathbb{Z}$: Indeed, this can be checked rationally, and rationally $g_{\ast}\mathbb{Q}$ is the free $E_{\infty}$-algebra on $\mathbb{Q}(-1)[-1]$. Thus, $g_{\ast}\mathbb{Z}$ is filtered by modules inflated along $g_{\ast}\mathbb{Z}\to\mathbb{Z}$, which implies descendability. ∎
@@
$\mathbb{G}_{m}^{\mathrm{rat}}=\mathrm{Spec}(\mathbb{Q}[T^{\pm 1}]\times_{\mathbb{Q}}\mathbb{Z}[\tfrac{1}{p}])$
(where the map $\mathbb{Q}[T^{\pm 1}]\to\mathbb{Q}$ is evaluation at $T=1$). We note that $\mathcal{D}_{\mathrm{qc}}(G_{k})$ is compactly generated by the powers $L^{\otimes n}$, $n\in\mathbb{Z}$. Indeed, to check that these dualizable objects are compact, it suffices to check that the unit is compact, for which it suffices to check that pushforward along $G_{k}\to\mathrm{Spec}(\mathbb{Z}[\tfrac{1}{p}])$ preserves colimits. This can be done fibrewise, but the fibres over $G_{k}$ are $\mathrm{Spec}(\mathbb{F}_{\ell})$ for finite $\ell\neq p$, and $\ast/\mathbb{G}_{m}$ over the generic point; in both cases, it is clear. Similarly, one proves that $L^{\otimes n}$ generate as they do in each fibre.
Note that $\mathcal{D}_{\mathrm{mot}}(k)$ is naturally linear over $\mathcal{D}_{\mathrm{qc}}(G_{k})$. Indeed, the map to $\ast/\mathbb{G}_{m}$ is given by the Tate twist, while the lift to the torsor over it is given by the isomorphisms
@@
The functor $\mathcal{D}_{\mathrm{qc}}(G_{k})\to\mathcal{D}_{\mathrm{mot}}(k)$ is fully faithful, with essential image being the subcategory generated under colimits, shifts, and retracts by $\mathbb{Z}(n)$ for $n\in\mathbb{Z}$. This agrees with $\mathcal{D}_{MT}(k)\subset\mathcal{D}_{\mathrm{mot}}(k)$.
@@
Recall that $\mathcal{D}_{\mathrm{qc}}(G_{k})$ is compactly generated by the tensor powers of the line bundle $L$. To prove the proposition, it remains to compare homomorphisms between generating objects, and this can be split up into the case of torsion coefficients, and rational coefficients. The case of modulo $n$ coefficients is clear as both categories become $\mathcal{D}(\mathbb{Z}/n\mathbb{Z})$, while the rational case follows from the previous proposition.
It is clear that the subcategory generated by all $\mathbb{Z}(n)$ is contained in $\mathcal{D}_{MT}(k)$. For the converse, pull back to $\mathcal{D}_{MT}(C)$ and use that every object is a retract of pull-push along $\mathcal{M}_{\mathrm{arc}}(C)\to\mathcal{M}_{\mathrm{arc}}(k)$. ∎
We note that it follows that $\mathcal{D}_{MT}(k)$ has a natural $t$-structure coming from the equivalence with $\mathcal{D}_{\mathrm{qc}}(G_{k})$. This is the “motivic $t$-structure”, whose connective part is the one generated under colimits by $\mathbb{Z}(n)$, $n\in\mathbb{Z}$; after $\ell$-adic completion, it gives the usual $t$-structure on $\ell$-adic sheaves.
@@
As before, $C$ denotes some completed algebraic closure of $k((T))_{1/2}$.
@@
induced by the map $f_{\ast}\mathbb{Z}\cong g_{\ast}\mathbb{Z}\to\mathbb{Z}$. It suffices to see that $f_{\ast}\mathbb{Z}\to\mathbb{Z}$ is coconnectively faithfully flat in $\mathcal{D}_{MT}(k)\cong\mathcal{D}_{\mathrm{qc}}(G_{k})$, i.e. tensoring preserves coconnective modules, and is conservative on them. Covering $G_{k}$ by $\mathrm{Spec}(\mathbb{Z}[\tfrac{1}{p}])$, we have to see that the map from the free $E_{\infty}$-$\mathbb{Z}[\tfrac{1}{p}]$-algebra on $\mathbb{Q}[-1]$ back to $\mathbb{Z}[\tfrac{1}{p}]$ is coconnectively faithfully flat. This is a simple computation. ∎
Let $G_{C}\to G_{k}$ be the algebraic stack parametrizing, in addition to the line bundle $L/R$ with compatible isomorphisms $L/n\cong\mu_{n}(k)\otimes_{\mathbb{Z}}R$, a map
@@
in $\mathcal{D}(\mathbb{Z})$ that after reduction modulo $n$ yields the natural map
@@
in $\mathcal{D}(\mathbb{Z})$ which after pullback to $\mathcal{O}_{C}^{\times}/(1+C_{<1})\otimes_{\mathbb{Z}}R\cong L\otimes\mathbb{Q}/\mathbb{Z}[\tfrac{1}{p}]$ is identified with
@@
As $C^{\times}/\mathcal{O}_{C}^{\times}\cong\mathbb{Q}$, the map $G_{C}\to G_{k}$ is a torsor under $\mathrm{Hom}(\mathbb{Q},L)$. This is an affine group scheme, and in fact $h:G_{C}\to G_{k}$ is a relative affine stack, so
$\mathcal{D}_{\mathrm{qc}}(G_{C})\cong\mathrm{Mod}_{h_{\ast}\mathcal{O}}\mathcal{D}_{\mathrm{qc}}(G_{k}).$
Moreover, $h_{\ast}\mathcal{O}$ is a complex which in degree 0 is given by $\mathcal{O}$, and in degree $1$ by $L^{\vee}\otimes\mathbb{Q}$, and 0 otherwise.
The identification $\mathbb{Z}(1)(C)\cong C^{\times}/(1+C_{<1})[-1]$ yields a natural $\mathcal{D}_{\mathrm{qc}}(G_{k})\cong\mathcal{D}_{MT}(k)$-linear symmetric monoidal functor
$\mathcal{D}_{\mathrm{qc}}(G_{C})\to\mathcal{D}_{MT}(C).$
@@
The functor $\mathcal{D}_{\mathrm{qc}}(G_{C})\to\mathcal{D}_{\mathrm{MT}}(C)$ is an equivalence. Consequently,
$\mathcal{D}_{\mathrm{mot}}(C)\cong\mathcal{D}_{\mathrm{mot}}(k)\otimes_{\mathcal{D}_{\mathrm{qc}}(G_{k})}\mathcal{D}_{\mathrm{qc}}(G_{C}).$
@@
By our preparations, it suffices to see that $h_{\ast}\mathcal{O}\cong f_{\ast}\mathbb{Z}$ as $E_{\infty}$-algebras under the equivalence $\mathcal{D}_{\mathrm{qc}}(G_{k})\cong\mathcal{D}_{MT}(k)$. The constructions give a natural map, and we know both sides explicitly as complexes: They are $\mathbb{Z}[\tfrac{1}{p}]\oplus\mathbb{Q}(-1)[-1]$. One only has to see that the map $\mathbb{Q}(-1)[-1]\to\mathbb{Q}(-1)[-1]$ is an isomorphism. But such endomorphisms are given by multiplication by some rational number, and it is easy to see that it must be nonzero. More precisely, the $\mathbb{Q}$ appearing is in both cases more canonically the value group of $C$, and if one uses this canonical description, the map is the identity. ∎
@@
and consequently a section of $G_{C}\to G_{k}$; this is responsible for the nearby cycles functor